Actions
Ops/Comm Scheduling meeting¶
Location : CR11 in R3 @ 15:30 pm on Tuesday 24th
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Agenda¶
Actions from last meeting- Generate appropriate table to store absorber pressure and material information (Janusz) : (almost done)
- DS readback current requires calibration (DL) : stands until close to powerup
- Henry, Craig, Paul and Pierrick to look into the water detectors in the trench : Still being worked on. Stands
- DC report
- Work plan for the next week
Ongoing work
- SS QD/QP commissioning
- Electrical work : http://micewww.pp.rl.ac.uk/attachments/6181/MICE_Electrical_Work_List_25APR16.docx
- Water system work
- FC cooldown
- Target fibre work
- Planning for July running
- A prototype run plan has been proposed (this is subject to change)
- First week of next user cycle (28th June - 4th July) reserved for bringing the magnet chain up (without M2 in SSD)
- We provide one reserve shifter a day during this period
- Rest of the user cycle (5th - 29th July) focussed on channel characterisation and zero absorber running with field. We may need to
run 24 hours a day for this period to take all the requested data in the time available.
- Shift calendar will be released by the end of this week or start of next week
- Should go through the Run Preparation List at next weeks meeting
- Update to work on alarm handler and spreadsheets
- A prototype run plan has been proposed (this is subject to change)
- AOB
- DONM : 31/3/2016 IN CR21 in R1
Minutes¶
- *Present : * Victoria B, Henry, Mark, Vicky B, Andy, Josef, Pierrick, Craig, John, Alan
- Actions from last meeting
- Table installed to store absorber pressure and material information but needs more info (Janusz) : (almost done)
- DS readback current requires calibration (DL) : stands until close to powerup
- Henry, Craig, Paul and Pierrick to look into the water detectors in the trench : Still being worked on. Stands
- Issues arising from last data taking run need to be identified and actioned (Boyd) : Stands (but I have a short list)
- DC report :
- FC cooldown started last Thursday, cooling trajectory looks good. Presently 228K
- Dump circuit racks using diodes good for 220C
- Diode and resistor packs for SSU & SSD now mounted in hall
- Primary School tour went well. 6 groups of ~10 students shown around
- Chiller concrete stand in place, cabling up presently
- R9 absorber moved and stand organised
- UPS units serviced in RR2 and H2 rooms
- PRY walls all in placed and closed and searched
- Hall water circuit for conventional magnets in progress - 1 week more
- TOF1&2, KL and EMR detectors surveyed
- PH made improvements to alarm handler now incorporating state machine
- Work plan for the next week : see attached
- Action Item Check ISIS spreadsheet (Boyd)
Updated by MacWaters, Craig about 7 years ago ยท 8 revisions